In a Poisson distribution, μ = 0.94. (Round the final answers to 4 decimal places.)

a. What is the probability that x = 0?

Probability          

b. What is the probability that x > 0?

Probability          
Answer #1

a) P(x = 0)

= 0.3906

b) P(x > 0)

= 1 - P(x = 0)

= 1 - 0.3906

= 0.6094